OpenStudy (anonymous):

PLEASE HELP! LAST PROBLEM! :D A straight road is to be constructed joining two towns A and B. If B is located 44 miles to the west and 30 miles to the south of A, find the length of the road and its bearing from A to B. The bearing from A to B is a. S 43 degrees W b. S 47 degrees W c. S 56 degrees W d. S 34 degrees W

OpenStudy (campbell_st):

draw a diagram  to find the length of the road use pythagoras' theorem \[x^2 = 44^2 + 30^2\] solve for x find the angle theta it will be the size of the angle south of west \[\tan(\theta) = \frac{30}{44}\] solve for theta \[\theta = \tan^{-1}\frac{30}{44}\] hope this helps